OpenStudy (anonymous):

can someone please just tell me what the answer to 5/6 -1/4 is i am gettin 1/6 but my answer key tells me a diiferent answer??

OpenStudy (anonymous):

alrigt, first get the same denominators:\[\frac{20 - 6}{24} = \frac{14}{24} = \frac{7}{12}\] ^_^

OpenStudy (anonymous):

do you understand it now? :)

OpenStudy (anonymous):

well i got 24 as the same denominator and then i just subtracted 5 from 1 and i guess thats where i went wrong right??

OpenStudy (anonymous):

yep ^_^ when you multiply to get the same denominator, you have to make sure to multiply both the numerators too by the number you've multiplied in the denomintor :)

OpenStudy (anonymous):

ohh i get it ok since i multiplied 6 times 4 and got 24 i have to multiply 5 times 4 as well?? then for the next one i did 486 and got 24 so i have to do 1* 6 is that right thats how u come up with 20/24-6/24=14/24 then the lowest form is 7/12

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

well done ^_^

OpenStudy (anonymous):

thank u a million times!!!!!!!!!!!!!

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

i have another question can u help me out again??

OpenStudy (anonymous):

the problem 15 3/5 + 9 1/2 i got 24 51/10 then down to simplest form 29 1/10 but my answer key says 25 1/10 can u tell me where i went wrong??

OpenStudy (anonymous):

You add the whole numbers by themselves first. So 15+9=24. 3/5+1/2 you need a common denominator which would be 10. So you have 6/10+5/10 which gives 11/10. 10/10 is equal to 1, so that can be added onto the whole number portion, so you have 24+1=25. Since you took 1 whole out of 11/10, you have 1/10 left. Leaving you with 25 1/10